Solve the system of equations by using subsitution. y = 5x + 6 and y = -2x - 8
Answer by mbmodak(32)

y=5x+6 (1)
y=-2x-8 (2) substitue for y from (2) in (1)
-2x-8=5x+6
-8-6=5x+2x
-14=7x
x=-(14/7)
x=-2 substitue for x in (1)
y=5*-2+6
y=-10+6
y=-4